How fast should a wire 20 cm long be moved through a 0.05-T magnetic field for an emf of 1 V to appear across its ends? 


1
Expert's answer
2021-11-07T19:58:31-0500

"E=Blvsin\theta"

"\theta=90°\\l=0.20m \\B=0.05T\\E=1V"

Put value

"v=\frac{E}{Blsin\theta}"

"v=\frac{1}{0.05\times0.2\times sin90°}"

"v=\frac{1000}{10}=100m/sec"